OpenStudy (anonymous):

if you re driving in 110 km/h along a straight road and you look to the side for 2.0 seconds, how long is your inattentive period?

OpenStudy (anonymous):

Speed = 110 km/h = 110 x 1000/3600 m/s = 275/9 m/s Time = 2.0 s Distance = speed x time = (275/9) x 2.0 = 61.1 m Ans: 61.1 m